OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0).

OpenStudy (anonymous):

y − 4 = −one half(x + 6) y − 4 = 2(x + 6) y + 6 = −one half(x − 4) y + 6 = 2(x − 4)

OpenStudy (anonymous):

we need to find the slope first by using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (-6,4) x1 = -6 and y1 = 4 (2,0) x2 = 2 and y2 = 0 now lets sub slope(m) = (0 - 4) / (2 - (-6) m = -4/8 = -1/2 <-- slope is -1/2 now we will use this formula : y - y1 = m(x - x1) slope(m) = -1/2 (-6,4) x1 = -6 and y1 = 4 now we sub y - 4 = -1/2(x - (-6) y - 4 = -1/2(x + 6) -- point slope form any questions ?